SOlve the linear system using elimination. Please show all work.
x+y=7 equation 1
y=-2x=8 equation 2

Answer by sheraj(2)

x+y=7
x=7-y
y-2x=8
y=8+2x
y=8+2(7-y)
y=8+14-2y
y=22-2y
3y=22
y=22/3
x=7-22/3=21-22/3
x=-1/3